# Rethinking Graph Backdoor Attacks: A Distribution-Preserving Perspective

### v1

@@ Appendix E EXPERIMENTS ON ATTACK TRANSFERABILITY
To demonstrate the transferability of our trigger generator in attacking various GNN architectures, we employ GCN as the surrogate model and evaluate the ASR and clean accuracy when attacking GCN, GraphSage (Hamilton et al., 2018) and GAT (Veličković et al., 2018), respectively. The results are presented in Tab. 4 – 6. From the tables, we observe that our DPGBA consistently achieves a high attack success rate while maintaining the clean accuracy across different target models and various defense settings. This indicates the adaptability and transferability of our framework, enhancing its practical value in real-world applications.
@@ Appendix F AGAINST VARIOUS OUTLIER DETECTION METHODS
To further demonstrate the in-distribution property of the triggers generated by our DPGBA, we adopt various state-of-the-art graph outlier detection methods, including DOMINANT (Ding et al., 2019), DONE (Bandyopadhyay et al., 2018) and its variant AdONE, AnomalyDAE (Fan et al., 2020), GAAN (Chen et al., 2020) and CONAD (Xu et al., 2022b), as defense mechanisms and conduct backdoor attacks on four datasets. The other settings are the same as the evaluation protocol in Sec. 5.1.3. The results of ASR are reported in Tab. LABEL:againstvarioustable. From the table, we observe that DPGBA consistently exhibits its capability to evade various graph outlier detection methods and maintain a high attack success rate. This consistency underscores the practical application value of DPGBA in real-world scenarios.
@@ Appendix G ADDITIONAL EXPERIMENTS
Datasets Defense Clean Graph UGBA Ours
@@ Appendix H DETAILS OF COMPARED METHODS
  • SBA-Samp (Zhang et al., 2021a): This method introduces a static subgraph as a trigger into the training graph for each poisoned node. The subgraph’s connections are formed based on the Erdos-Renyi (ER) model, while its node features are randomly selected from those in the training graph.
@@ Appendix I IMPLEMENTATION DETAILS
During the bi-level optimization phase, the computation cost of each outter iteration consist of updating of surrogate GCN model and OOD detector in innter iterations and training adaptive trigger generator. Let $h$ denote the embedding dimension. The cost for updating the surrogate model is approximately $O(Nhd|\mathcal{V}|)$, where $d$ is the average degree of nodes and $N$ is the number of inner iterations for the surrogate model, which is generally small. The cost for updating the OOD detector is approximately $O(Khd|\mathcal{V}|)$, where $K$ is the number of inner iterations for the OOD detector. For trigger generator, the cost for optimizing $\mathcal{L}_{T}$ is $O(hd|\mathcal{V}|)$, for optimizing $\mathcal{L}_{D}$ is $O(hd|\mathcal{V}|)$, and for optimizing $\mathcal{L}_{E}$ is $O(hd|\mathcal{V}|+|\mathcal{V}|^{2}h+|\mathcal{V}||\mathcal{V}_{t}||\mathcal{V}_{L}|h)$, where $|\mathcal{V}_{L}|$ and $|\mathcal{V}_{t}|$ are generally small compared to $|\mathcal{V}|$. In the test phase, the cost of generating trigger to attack the target node is only $O(h)$. In our empirical experiments conducted on large-scale datasets, such as Flickr and OGB-arxiv, which comprise 899,756 and 169,343 nodes respectively, we streamlined the training process by selecting a subset of $\mathcal{V}$ and setting $|\mathcal{V}|=4096$ for each epoch. Despite this simplification, DPGBA still achieve a high attack success rate, as evidenced in Tab. 4 – 6. These results demonstrate that the DPGBA effectively generates triggers that the victim model quickly memorizes, underscoring its great potential for conducting scalable target attacks.

### v2

@@ Appendix E EXPERIMENTS ON ATTACK TRANSFERABILITY
To demonstrate the transferability of our trigger generator in attacking various GNN architectures, we employ GCN as the surrogate model and evaluate the ASR and clean accuracy when attacking GCN, GraphSage (Hamilton et al., 2017) and GAT (Veličković et al., 2018), respectively. The results are presented in Tab. 4 – 6. From the tables, we observe that our DPGBA consistently achieves a high attack success rate while maintaining the clean accuracy across different target models and various defense settings. This indicates the adaptability and transferability of our framework, enhancing its practical value in real-world applications.
@@ Appendix F AGAINST VARIOUS OUTLIER DETECTION METHODS
To further demonstrate the in-distribution property of the triggers generated by our DPGBA, we adopt various state-of-the-art graph outlier detection methods, including DOMINANT (Ding et al., 2019), DONE (Bandyopadhyay et al., 2019) and its variant AdONE, AnomalyDAE (Fan et al., 2020), GAAN (Chen et al., 2020) and CONAD (Xu et al., 2022b), as defense mechanisms and conduct backdoor attacks on four datasets. The other settings are the same as the evaluation protocol in Sec. 5.1.3. The results of ASR are reported in Tab. LABEL:againstvarioustable. From the table, we observe that DPGBA consistently exhibits its capability to evade various graph outlier detection methods and maintain a high attack success rate. This consistency underscores the practical application value of DPGBA in real-world scenarios.
@@ Appendix G ADDITIONAL EXPERIMENTS
Datasets Defense Clean Graph UGBA DPGBA
@@ Appendix H DETAILS OF COMPARED METHODS
  • SBA-Samp (Zhang et al., 2021): This method introduces a static subgraph as a trigger into the training graph for each poisoned node. The subgraph’s connections are formed based on the Erdos-Renyi (ER) model, while its node features are randomly selected from those in the training graph.
@@ Appendix I IMPLEMENTATION DETAILS
During the bi-level optimization phase, the computation cost of each outter iteration consist of updating of surrogate GCN model and OOD detector in inner iterations and training adaptive trigger generator. Let $h$ denote the embedding dimension. The cost for updating the surrogate model is approximately $O(Nhd|\mathcal{V}|)$, where $d$ is the average degree of nodes and $N$ is the number of inner iterations for the surrogate model, which is generally small. The cost for updating the OOD detector is approximately $O(Khd(|\mathcal{V}_{s}|+|\mathcal{T}_{P}|))$, where $K$ is the number of inner iterations for the OOD detector. For trigger generator, the cost for optimizing $\mathcal{L}_{T}$ is $O(hd|\mathcal{V}_{U}|)$, for optimizing $\mathcal{L}_{D}$ is $O(hd(|\mathcal{V}_{s}|+|\mathcal{T}_{P}|))$, and for optimizing $\mathcal{L}_{E}$ is $O(hd|\mathcal{V}|+|\mathcal{V}_{U}|^{2}h+|\mathcal{V}_{U}||\mathcal{V}_{t}||\mathcal{V}_{L}|h)$, where $|\mathcal{V}_{L}|$ and $|\mathcal{V}_{t}|$ are generally small compared to $|\mathcal{V}|$. In our empirical experiments conducted on large-scale datasets, such as Flickr and OGB-arxiv, which comprise 899,756 and 169,343 nodes respectively, we streamlined the training process by selecting a subset of $\mathcal{V}_{U}$ and setting $|\mathcal{V}_{U}|=4096$ for each epoch. Despite this simplification, DPGBA still achieve a high attack success rate, as evidenced in Tab. 4 – 6. In Table 9, we report the overall training time and corresponding ASR of our DPGBA compared to GTA and UGBA on the OGB-arxiv dataset. All models were trained on a Nvidia A6000 GPU with 48GB of memory. The results indicate that DPGBA requires only approximately 20 seconds more training time compared to the baselines on the OGB-arxiv dataset. Given that our DPGBA achieves an ASR of over 90%, while the baseline methods achieve nearly 0% with OD defense adopted, this additional time is justified. This demonstrates that DPGBA effectively generates triggers that the victim model quickly memorizes, highlighting its potential for conducting scalable targeted attacks.
Table 9. Training Time
Metrics GTA UGBA DPGBA
ASR (None) 92.0 94.3 92.7
ASR (OD) 0.00 0.00 92.0
Time 37.7s 41.8s 60.6s